OpenStudy (anonymous):

A simple quadratic equation test :) Medal for the person you reply first and "CORRECT"

OpenStudy (anonymous):

Solve 6x2 + 11x – 35 = 0

OpenStudy (anonymous):

\[-\frac{ 7 }{ 2 } , \frac{ 5 }{ 3 }\] are the two root. Now solve it :)

OpenStudy (anonymous):

\[ 6 x^2+11 x-35=(2 x+7) (3 x-5) =0\\ x= -\frac 7 2\\ x= \frac 5 3 \]

OpenStudy (anonymous):

@eliassaab sir you are really fast :D